A 90 kg man runs at 5 m/s. What is his momentum (p)?




Vector A-(1,3) m and vector B (-1,-3) m. Find C D A-B. Also find the magnitude and direction of D. A+B and
Answer #1

1.

Momentum = Mass*Velocity

P = m*V

P = 90*5 = 450 kg-m/sec

2.

A = (1, 3) = i + 3j

B = (-1, -3) = -i - 3j

C = A + B

C = (i + 3j) + (-i - 3j) = 0 i + 0 j

C = (0, 0) m

D = A - B

D = (i + 3j) - (-i - 3j) = i + 3j + i + 3j

D = 2i + 6j = (2, 6) m

|D| = sqrt (2^2 + 6^2) = 6.32 m

Direction = arctan (6/2) = 71.56 deg Counter-Clockwise to +x-axis

Please Upvote.